Q–2: [5+2+3 Marks] Let X be a random variable giving the number of
heads minus the number of tails in three tosses of a coin.
a) Find the probability distribution function of the random variable X.
b) Find P(−1 ≤ X ≤ 3).
c) Find E(X).
Answer #1


Formulas: 1) If P(X-x) is probability distribution function of the discrete random variable X then E(X) xx * P(X-x). Given th Number of possible ways to get 0 heads and 3 tails [TTT] 1 P(X =-3) = P(0 heads and 3 tails) = 8 Number of possible ways to ga) Now we have to find the probability distribution function of the random variable X- P(X -x) -1 P(X = x) 8 b) Now we have tc) Now we have to find E(X) (-3 * P(X =-з)) + (-1 * P( -1)) + (1 * P(X = 1)) + (3 * P(X = 3)) 0